Karla:

An object falls from rest. In the last second of its fall it covers half of the total distance. If g=9.8m/s^-2 then the total time of its fall is A)(2-2^1/2)s B)(2+2^1/2)s C)(2+-2^1/2)s D)4s

sillybilly123:

Using a SUVAT eqn, displacement \(s(t) = ut + \frac{1}{2} a t^2\), BUT with \(u = 0\) as object falls from rest. So: \(s(t) = \frac{1}{2} a t^2\). As \(a = g\), with downward velocity positive, if the total travel time is \(T\) then: \(\qquad s(T) = \frac{1}{2} g T^2\) Equally, distance travelled up to "start of last second" is: \(\qquad s(T-1) = \frac{1}{2} g (T-1)^2\) So, distance travelled in last second is: \(\qquad s(T) - s(T-1) = \frac{1}{2} g \left( T^2- (T-1)^2 \right)\) \(\qquad = \frac{1}{2} g \left( 2T - 1 \right)\) Using given ratio: \(\qquad \dfrac{\frac{1}{2} g \left( 2T - 1 \right)}{ \frac{1}{2} g T^2} = \dfrac{1}{2}\) \(\qquad \implies T^2 - 4T + 2 = 0\) Leaving you to solve that (using the quadratic formula) to get solution C)

Karla:

Thank you!

Zarkon:

An interesting way to do this problem is to use the conservation of mechanical energy

Karla:

Oh, okay Thank you!